OpenStudy (anonymous):

If p equals two , what is the value of three p plus two minus p raised to the second power ? Show your work.

jimthompson5910 (jim_thompson5910):

p = 2 3p+2-p^2 3(2)+2-(2)^2 I'll let you finish

OpenStudy (anonymous):

I'm confused at the end.

jimthompson5910 (jim_thompson5910):

3*2 = ?

OpenStudy (anonymous):

6

jimthompson5910 (jim_thompson5910):

2^2 = 2*2 = ?

OpenStudy (anonymous):

4

jimthompson5910 (jim_thompson5910):

3(2)+2-(2)^2 then becomes 6+2-4

OpenStudy (anonymous):

4

jimthompson5910 (jim_thompson5910):

exactly

OpenStudy (anonymous):

so the value is 4?

jimthompson5910 (jim_thompson5910):

3p+2-p^2 = 4 when p = 2